The length of a rectangle is 6 in. more than twice its width. If the perimeter of the rectangle is 30 in., find the width of the rectangle
Answer by Earlsdon(6294)

The perimeter of a rectangle is given by:
P = 2L + 2W Substitute: P = 30, L = 2W+6, and W = W
30 = 2(2W+6) + 2W Simplify.
30 = 4W+12+2W
30 = 6W+12 Subtract 12 from both sides.
18 = 6W Divide both sides by 6
3 = W
The width of the rectangle is 3 inches.